I must solve for y in the equation: y/5-7=-2x. What does y equal.
Found 2 solutions by Alan3354, midwood_trail:
Answer by Alan3354(69443)

I must solve for y in the equation: y/5-7=-2x. What does y equal.
------------
y/5-7=-2x
Multiply by 5
y-35 = -10x
Add 35
y = 35-10x



Answer by midwood_trail(310)

Here is my reply to another question very similar to this one.
Follow the steps below and you will be able to solve for y in this case.
To solve for a variable means to isolate the variable.
In other words, we need to place x all alone on one side of the equation.
-3x + 7 = 2y + 3
Subtract 7 from both sides.
-3x = 2y + 3 - 7
-3x = 2y - 4
We now divide both sides by -3 to isolate x.
x = (2y - 4)/-3
Did you follow?